OpenStudy (underreconstruction):

What is the value g(-3) when g(x) = 2x + 1? 7 -5 5 -6

OpenStudy (mathmate):

Hint and example if f(x)=2x+1, then f(5)=2(5)+1=11.

ganeshie8 (ganeshie8):

That means you simply have to replace `x` by `-3` and simplify

OpenStudy (daddysbabygirl):

Do you still need help with this ?

OpenStudy (okkaaayyyy):

The correct answer is 5. To get this you substitute x for -3 in this equation.

OpenStudy (mathmate):

@Underreconstruction g(-3)=2(-3)+1=?

OpenStudy (underreconstruction):

g=5/3

OpenStudy (brainzonly):

What is the value g(-3) when g(x) = 2x + 1 g(-3) = 2(-3) + 1 -6+1 g(-3) = -5 B

RhondaSommer (rhondasommer):

sub x for -3. 2(-3)+1 -6+1 -5 "_" hope that helps.

OpenStudy (underreconstruction):

Thank you! :D